OpenStudy (anonymous):

Trying to find displacement and total distance traveled. a(t) = 13; v-sub(zero) = -1; 0 < t <2

OpenStudy (anonymous):

\(a(t)=13\) right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So I get that v(t) = 13t + C and C = -1 at t = 0

OpenStudy (anonymous):

pull back one level to get \(v(t)=13t+C\) and since \(v(0)=-1\) it must be \(v(t)=13t-1\)

OpenStudy (anonymous):

Then I should go one step further to get s(t)? Then I get 13t^2 - t + C

OpenStudy (anonymous):

it is not "C = -1 at t = 0 " but rather \(C=-1\) period

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

now like before you have to integrate right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

That's when I get \[13t ^{2} - t + C\]

OpenStudy (anonymous):

not "take the anti derivative" although you will need that

OpenStudy (anonymous):

but this time it is a definite integral, not an indefinite one with that \(+C\) out at the end

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

good question

OpenStudy (anonymous):

you want an actual number out of this right? not just some function whose derivative is \(13x-1\)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

in other words, you want \[\int_0^213t-1dt\]

OpenStudy (anonymous):

the way you do this is by finding an anti derivative of \(13x-1\) namely \(\frac{13}{2}x^2-x\) evaluate at the upper limit, the lower limit, and subtract or you could just use plain geometry because all you are looking at is a line

OpenStudy (anonymous):

I failed to see the difference between indefinite and definite integration I think.

OpenStudy (anonymous):

but if you have a second, we can go back to the original question, why not the \(+C\) and see if i can explain clearly

OpenStudy (anonymous):

we could do it in a general case, or work with a specific example

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

ok lets do it with a specific example, lets let \(f(x)=2x+1\)

OpenStudy (anonymous):

notice i wrote \(f(x)\) because i am thinking that \(f\) is a function of \(x\)

OpenStudy (anonymous):

there are an infinite number of "anti derivatives" for \(f\), anything that looks like \[F(x)=x^2+x+C\] but they all have to look like that, because if two functions have the same derivative, they can only differ by a constant i.e. it has to be \(x^2+x+number\)

OpenStudy (anonymous):

Gotcha

OpenStudy (anonymous):

now lets consider this specific function of \(x\) \[A(x)=\int_2^x(2t+1)dt\] notice that it is a function of \(x\) and not of \(t\)

OpenStudy (anonymous):

i.e. \[A(1)=\int_2^1(2t+1)dt\] and \[A(3)=\int_2^3(2t+1)dt\] etc

OpenStudy (anonymous):

we know from the fundamental theorem of calculus that \[A'(x)=f(x)=2x+1\]

OpenStudy (anonymous):

which means what do we know about \(A\)? it must be \[A(x)=\int_2^x(2t+1)=x^2+x+C\] but \(A\) is not some general function, it is a specific function, so \(C\) is not some general constant, it is a fixed number we know

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

and it is easy to find, because we know one point on the graph of \(A\) namely we know that \[A(2)=\int_2^2(2t+1)dt=0=2^2+2+C\] telling us that \(C=-(2^2+2)\)

OpenStudy (anonymous):

that is why we can write \[A(x)=\int _2^x(2t+1)dt=x^2+x-6\] a very specific function, and also why we can compute \[A(5)=\int_2^5(2t+1)dt\] by taking an anti derivative of \(2x+1\) which is \(F(x)=x^2+x\) evaluate at the upper limit, evaluate at the lower limit, and subtract when you evaluate \(F(x)=x^2+x\) at \(2\) you get the \(-6\) that you need to subtract off, i.e. you get \(C=-6\)

OpenStudy (anonymous):

well that was rather long winded, but i hope it made some sense

OpenStudy (anonymous):

It helped make sense of what was explained in class. It's good to see things from other people

OpenStudy (anonymous):

ok good now we still didn't solve your problem, the displacement and the total distance, but in any case you should be able to do it now btw you got the anti derivative wrong, it should be \(\frac{13}{2}x-x\)

OpenStudy (anonymous):

Yeah. I realized that. I do that a lot it seems. But I was able to figure out the answers, thanks again!

OpenStudy (anonymous):

yw